

********************************************************************
UARS WEEKLY STATUS REPORT - 29 Oct 2004
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM successfully gathered daily mid (1.1 nm) resolution, MgII,
occultation, solar pointing offset, and continuous monitoring
scans for time periods with UARS telemetry or tape recording.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004189 has been reprocessed and archived at the GSFC DAAC.
MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 2004159. MEPS data files from 2004160
to 2004179 are processed and waiting for archiving.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing recieved.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 16 October 2004 (Orbit 71676, GMT Day 290) through
22 October 2004 (Orbit 71781, GMT Day 296).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 292/2315 71721 AXIS 1 & 2 OFF
PEM 292/2341 71721 ZEPS / NEPS DAY ONLY
HALOE 293/1330 71730 POWER ON
The beta angle ranged between the angles of 63.3 degrees to 42.4
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 02 November 2004 (DOY 307).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.63 to 0.97 degrees C. Additional
battery performance data for each battery on 15 October and 22 October
2004 are provided in Appendix D.
The clock error ranged from 10.1 to -10.9 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 05
PEM = 02
HALOE = 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 06
Four (4) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Six (6) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the FOT attended a meeting with FDF and CMS software
development to discuss the issues surrounding the use of the SMA
service on TDW. Currently, FDF delivers a User Antenna View (UAV)
file to the User Planning System (UPS) and Command Management System
(CMS). This file presently contains only SSA and MA view period
information for TDRSs East, West and 275, but does not contain the
SMA view period information needed to utilize the automatic "Recurrent
Scheduling" feature in the UPS. Consequently, the FOT has been unable
to automatically schedule SMA TDRS events during forecast scheduling.
Therefore, the FOT met with FDF to discuss the ramifications of adding
SMA view period data to the UAV file. However, since the UAV file is
delivered to the CMS, as well as the UPS, any changes made to the file
to facilitate UPS scheduling requirements might necessitate a software
change to the CMS. As it turns out, there is a processing option in
the UPS that utilizes MA view periods in the UAV file to formulate SMA
view periods. (This option is a valid work-around because times for
SMA and MA view periods are identical.) After discovering this
processing option, the FOT enabled it and performed a test run of it's
forecast scheduling process. A few "test" SMA events were
automatically generated on TDW using the Recurrent Scheduling feature.
These test events were transmitted to White Sands Scheduling as part
of the weekly forecast batch on October 19, 2004. The FOT will let
these SMA events flow through the normal forecast scheduling process.
As soon as the events are confirmed by White Sands Scheduling, they
will be deleted by the FOT, thus closing the loop on the SMA view
period work-around. If the work-around is successful, no
modifications to the FDF- provided UAV file will be necessary.
MEETINGS
-None this report period
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 02 November 2004 (DOY 307).
This maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 22 October 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
71718 292/1915 04-297 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
71721 292/2315 04-298 PEM AXIS 1 & 2 OFF
71721 292/2341 04-299 PEM ZEPS/NEPS DAY ONLY
71730 293/1330 04-300 HALOE POWER ON
71731 293/1514 04-301 HALOE UPDATE DAC VALUES
71761 295/1508 04-302 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
71776 296/1549 04-303 CDH DSN XMTR CMDS
71777 296/1645 04-304 CDH DSN XMTR CMDS
71780 296/2153 04-305 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 22 October 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 22 October 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
71718 292/1918 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
71761 295/1510 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
71780 296/2158 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 22 October 2004
15 October 2004 - GMT Day 289 Beta = 64.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.68/-4.94 +00.0/+00.0 07.0 0.00 0.0
2 -28.0/- 78.4 +3.87/+2.92 +20.0/-14.4 26.7 1.04 16.7
3 -33.6/-179.2 +2.60/+1.66 +18.8/-13.2 26.7 1.04 15.7
22 October 2004 - GMT Day 296 Beta = 42.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.31 +00.0/+00.0 07.0 0.00 0.0
2 11.2/- 50.4 +6.11/+5.14 +25.6/-15.6 26.2 1.04 19.6
3 -11.2/-168.0 +4.50/+3.87 +24.0/-15.2 26.2 1.03 18.8
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period
APPENDIX E
UARS ANOMALY REPORTS
Week ending 22 October 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-073 FOT 71691 291/0026 37270 Data Drop Outs
04-066 FOT 71717 292/1623 30832 Generic Late Acquisition
04-074 FOT 71730 293/1330 37291 30 sec Data Dropout
04-066 FOT 71746 294/1540 30832 Generic Late Acquisition
04-066 FOT 71770 296/0552 30832 Generic Late Acquisition
04-075 FOT 71776 296/1549 N/A Canberra Antenna Problem
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-068 FOT 71572 283/0233 37190 WSGT DROPOUT/FAILOVER
04-069 FOT 71587 284/0230 37194 WSGT Equip Failover
04-070 FOT 71590 284/0559 37195 WSGT Dropout
04-071 FOT 71616 286/0043 N/A PEM IN WRONG MODE - OFF!
04-072 FOT 71648 288/0302 N/A SSPP In stow due to eclipse
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-066 FOT See List Oct 04 30832 Generic Late Acquisition
04-073 FOT 71691 291/0026 37270 Data Drop Outs
04-074 FOT 71730 293/1330 37291 30 sec Data Dropout
04-075 FOT 71776 296/1549 N/A Canberra Antenna Problem